A 20kg body is allowed to roll do an inclined plane 4m long with an inclination 30 degree. Calculate the work done. (g=10m/s)
1
Expert's answer
2019-07-09T11:24:50-0400
"F = mg\sin\alpha"

"F = 20*10*0.5 = 100 N"

"A = FS""A = 100*4 = 400 J"

Answer: 400 J